Manufacturer made two announcements concerning its common stock today.


11. Manufacturer made two announcements concerning its common stock today. First, the company announced that the next annual dividend will be $1.75 a share. Secondly, all dividends after that will decrease by 1.5 percent annually, What is t maximum you should pay to purchase the stock if you require a 14% return? 

A. $11.29 B.$12.64 C.$13.27 D.$14.00 E.$14.21 


12. The common stock of Textile Mills pays an annual dividend of $1.65 a share. The Company has promised to maintain a constant dividend even though the economy is weak. How much should you pay for the stock to glean a 12% rate of return?

A. 12 13.75 B.14.01 C.14.56 D.14.79 E.15.23 


13. The price of common stock (Po) depends on (1) the cash flows investors expect to receive if they buy the stock and (2) the riskiness of the expected cash flows. 

A.True B.False 


14. The common stock of Auto Deliveries sells for $28.16 a share. The stock is expected to pay a $1.35 dividend next year and the market expects the firm to grow by 3% annually. What is the rate of return on the stock?

A.7.42% B.7.79% C.19.67% D.20.14% E.20.86% 


15. Northern Gas recently paid a $2.80 annual dividend on its common stock. This dividend increases at a rate of 38% per year. The stock is selling for $26.91. what 15 is the market expected rate of return? 

A. 13.88% B. 14.03% C. 14.21% D. 14.37% E. 14.60% 
Answer #2

SOLUTION :


11.


Given :


D1 = 1.75 ($)

g = - 1.5 % = - 0.015

r = 14 % = 0.14


Now,


Stock is  worth, P0

= D1 / (r - g)

= 1.75/(0.14 - (- 0.015))

= 11.29  ($)



So,  maximum price that can be paid = Option A. $11.29 (ANSWER). 


12. 


Given :


D0 = 1.65 ($) 

g = 0 % (constant dividend)

r = 12% = 0.12

So,


Stock is  worth, P0

= D1 / (r - g)

= 1.65/(0.12 - 0))

= 13.75  ($)



So,  maximum price that can be paid = Option A. $13.75 (ANSWER). 



13.


False  (since riskiness of cash flows is not considered in the dividend model)  (ANSWER) .


14.


Given :


P0 = 28.16 ($)

D1 = 1.35 ($)

g = 3 % = 0.03



Now,


r =  D1 / P0 + g

= 1.35/28.16 + 0.03

= 0.0779

= 7.79 % 


So, rate of return on the stock = Option B. 7.79 % .  (ANSWER).


15.


Given :


P0 = 26.91 ($)

D0 = 2.80 ($)

g = 3.8 % = 0.038



Now,


r =  D1 / P0 + g

= (2.80*(1 + 0.038)) / 26.91 + 0.038

= 0.1460

= 14.60% 


So, rate of return on the stock = Option E. 14.60 % .  (ANSWER).





answered by: Tulsiram Garg
Answer #1
11) D1 1.75
According to the dividend growth model.
P0 = D1/(R-g)
where R is .14
g -0.015
P0 (1.75/(.14-(-.015))
P0 11.29
A. $11.29.
12) D1 1.65
According to the dividend growth model.
when the dividend is constant
P0 = D1/(R)
where R is .12
P0 (1.65)/(.12)
P0 13.75
A. $13.75.
13) The price of a stock is given by the present value of
future dividend payments. The present value is calculated
based on an interest rate.
TRUE.
14) D1 1.35
According to the dividend growth model.
P0 = D1/(R-g)
where R is the rate of return.
g 0.03
P0 28.16
28.16 = 1.35/(R - .03)
(R-.03) = 1.35/(28.16)
(R-.03) = .04794
R = .04794 + .03
R = .07794
B. 7.79%.
15) D1 2.8
According to the dividend growth model.
P0 = D1/(R-g)
where R is the rate of return.
g 0.038
P0 26.91
26.91 = 2.8/(R - .038)
(R-.038) = 2.8/(26.91)
(R-.038) = .104051
R = .104051 + .038
R = .142051
C. 14.21%.